OpenStudy (anonymous):

solve the rational equation. check your answer. 1/2 + 3/t = 5/8

hero (hero):

\[\frac{1}{2} + \frac{3}{t} = \frac{5}{8}\]

hero (hero):

Subtract 1/2 from both sides to get: \[\frac{3}{t} = \frac{5}{8} - \frac{1}{2}\]

hero (hero):

The equivalent of 1/2 is 4/8 so: \[\frac{3}{t} = \frac{5}{8} - \frac{4}{8}\]

OpenStudy (anonymous):

I know my answer should be 24, but I don't know how to show my work.

hero (hero):

left side reduces to \[\frac{3}{t} = \frac{1}{8}\]

hero (hero):

I mean, right side, lol

hero (hero):

And then cross multiply to finish.

OpenStudy (anonymous):

so I just cross multiply the 3 and the 8?

OpenStudy (anonymous):

I also need to know how to check my answer.